A speedometer reads a speed at 103 km/hr. However, the actual speed the car is going is 98km/hr. What would be the percentage error?
Options are,

1)4.8
2)5.0
3)4.9
4)5.1
1
Expert's answer
2020-10-07T07:30:45-0400

The error of measurement of a speed

"\varepsilon=\frac{\Delta v}{v}\times 100\%\\
=\frac{103-98}{98}\times 100\%=5.1\%"

Answer: (4)